OpenStudy (anonymous):

The cost of coffee is determined by the type of coffee beans that are mixed together. The cost of a local mix of Arabica and Robusta beans is modeled by the equation below, where A = the pounds of Arabica and R = pounds of Robusta. : 700A + 1200R = 1,000,000 If a mixture contains 1000 pounds of Arabica beans, how many pounds of Robusta beans should be in the mixture? R = ________________ pounds

OpenStudy (anonymous):

700(1000)+1200R=1000000

OpenStudy (anonymous):

700000+1200R=1000000

OpenStudy (anonymous):

1200r=300000

OpenStudy (anonymous):

divide by 1200

OpenStudy (anonymous):

250 pounds

OpenStudy (anonymous):

Do you understand it how i got that formula and did the steps?

OpenStudy (anonymous):

700A+1200R=1,000,000 why the cost should be fixed at 1,000,000 is beyond me, but anyway that is what it says then it says A=1000 so now the equation is 700×1000+1200R=1,000,000 or 700,000+1200R=1,000,000 which i would solve in two steps 1) subtract 7000,000 to get 1200R=300,000 2) then divide by 1200 to get R=250